solve for r, 36r^2-1=0

Answer by Maths68(1474)

solve for r, 36r^2-1=0
36r^2-1=0
(6r)^2-(1)^2=0
(6r-1)(6r+1)=0
(6r-1)=0 or(6r+1)=0
6r=1 or6r=-1
r=1/6 or r=-1/6